# Budgeted Spatial Data Acquisition: When Coverage and Connectivity Matter

### v2

@@ 3.2 Problem Definition
Suppose we have a marketplace $\mathcal{D}=\{D_{1},\dots,D_{5}\}$, the corresponding dataset prices are $\{5,6,4,4,2\}$. Fig. 3(a) shows the corresponding cell-based datasets $\mathcal{S}_{\mathcal{D}}$ constructed form $\mathcal{D}$ and a grid $\mathcal{C}_{3,3}$. We then construct the spatial dataset graph $\mathcal{D}_{\mathcal{S_{D}},\delta}$ with $\delta=2$. For one of the dataset pairs $S_{D_{1}}$ and $S_{D_{2}}$, we can obtain the cell-based dataset distance by calculating the distance between their nearest neighbor cells, as shown in Fig. 3(b). Since the distance $dist(S_{D_{1}},S_{D_{2}})=1<\delta$, $S_{D_{1}}$ and $S_{D_{2}}$ are directly connected. After computing the distance between all pairs of datasets, we can obtain the spatial dataset graph, as shown in Fig. 3(c). A data buyer with a budget of \$15 wants to buy a set of connected datasets. From all feasible solutions of BMCC we can obtain the optimal subset =$\mathcal{H}^{*}=\{S_{D_{1}},S_{D_{2}},S_{D_{4}}\}$.
@@ 4 NP-hardness of BMCC
Given an arbitrary instance of the MCP, we sort the universe $\Omega$ according to the lexicographic order, then create a mapping from the sorted universe $\Omega=\{o_{1},o_{2},\dots,o_{|\Omega|}\}$ to an integer set $\{0,1,\dots,|\Omega|-1\}$. Then we can construct a $2^{\theta}\times 2^{\theta}$ grid in the space such that $2^{\theta}\times 2^{\theta}\geq|\Omega|$. Since each cell in the grid can be represented by a unique integer ID, each element in $\Omega$ can be mapped to a cell whose ID is equal to the integer, as shown in Figures 4(a) and 4(b). The other four cells with IDs 12, 13, 14, and 15 are not included in $\forall S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$, indicating that no datasets are located in those cells.
Then we set the price of each vertex to 1, the budget to $k$, and $\delta$ to $2^{\theta}\sqrt{2}$ in Definition 10. The set $\mathcal{S}_{\mathcal{D}}$ can be modeled into a complete undirected graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D},\delta}}=(\mathcal{V},\mathcal{E})$: each vertex $v_{i}$ corresponding to each cell-based dataset $S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$; for each vertex $v_{i}$ in the graph, there is an undirected edge $e_{i,j}$ $(\forall v_{j}\in\mathcal{V}$, $i\neq j)$. This is because $dist(S_{D_{i}},S_{D_{j}})$ must not exceed $\delta$, ensuring the connectivity constraint is always satisfied. Figure 4(c) shows a spatial dataset graph constructed from the MCP instance, and we can see that the total reduction is performed in polynomial time.
@@ 5 Approximation Algorithms for BMCC
In this section, we focus on solving the BMCC problem (defined in Def. 10). By Theorem 4.1, we know that this problem is NP-hard. Achieving the optimal solution for BMCC is computationally prohibitive. Therefore, we first propose a greedy algorithm, which provides a $\frac{p_{min}}{B}$ approximation guarantee (see Section 5.1). However, its approximation ratio decreases when confronted with a large budget $B$. To address this, we propose an improved greedy algorithm that produces a superior approximation ratio for scenarios involving large $B$ (see Section 5.2).
@@ 5.1 Dual-Search Algorithm (DSA)
In this subsection, we present a Dual-Search Algorithm called DSA. The basic idea of DSA is to perform a two-round heuristic search based on the greedy strategy. In the first-round search, the algorithm selects the dataset that maximizes the marginal gain, w.r.t. the ratio of the spatial coverage increment over the price of the dataset, while maintaining the spatial connectivity at each step. In the second-round search, it chooses the set that maximizes the marginal gain, w.r.t. the spatial coverage increment, while maintaining the spatial connectivity at each step. The optimal selections in the two-round search can result in two feasible solutions, which jointly guarantee the effective approximation ratio of DSA whenever the dataset chooses any pricing function. In the following, we will introduce how DSA works according to the pseudocode of Algorithm 1 and then theoretically prove the approximation ratio of DSA. Afterwards, we will further analyze its time complexity in detail.
@@ 25                   $\mathcal{U}_{2}=\mathcal{U}_{2}\backslash S_{D_{i}}$;
Fig. 5(a) shows a set of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\dots,S_{D_{8}}\}$, a distance threshold $\delta=1$ and a budget $B=\$6$. When performing DSA, we first find the candidates $\mathcal{S}_{\mathcal{D}}^{\prime}$. Since the price of each dataset $S_{D}\in\mathcal{S}_{\mathcal{D}}$ is less than $B$, the candidate datasets $\mathcal{S}_{\mathcal{D}}^{\prime}=\mathcal{S}_{\mathcal{D}}$ in this example. The first-round search preferentially selects the dataset with the maximum ratio $|S_{D_{i}}\cap\mathcal{U}_{1}|/{p(S_{D_{i}})}$ within the budget in each iteration. As shown in Fig. 5(b), the algorithm adds $\{S_{D_{8}},S_{D_{7}},S_{D_{4}}\}$ to the result subset $\mathcal{H}_{1}$ one by one until the budget is exhausted, where the spatial coverage of $\mathcal{H}_{1}$ is 8. Next, the second-round search iteratively selects the dataset with the maximum coverage increment $|S_{D_{i}}\cap\mathcal{U}|$ and satisfies spatial connectivity. Thus, we add the set $\{S_{D_{1}},S_{D_{5}}\}$ into $\mathcal{H}_{2}$, resulting in a spatial coverage of 10 for $\mathcal{H}_{2}$, as shown in Fig. 5(c). Finally, the algorithm returns the result $\mathcal{H}_{2}$, which contains a larger spatial coverage than $\mathcal{H}_{1}$.
@@
Due to the maximum spatial coverage of a set in the result subset is not less than the average spatial coverage of all sets, we have
@@ 38             $\mathcal{H}\leftarrow\mathcal{H}\cup\{S_{D_{i}}\}$;
In the next subsection, we will describe the details of DPSA. Essentially, we try to find a collection of paths with a common node in the spatial dataset graph (defined in Def. 9), where the path is represented as a sequence of nodes connected by edges. This ensures that the subset found is a connected subgraph. However, finding a combination of paths with a common node that maximizes the approximation ratio can be time-consuming (see the time complexity analysis in Section 5.2.1). To improve the efficiency, we propose two accelerate strategies that produce competitive solutions with DPSA in practice while achieving significant speedups in Section 5.2.2.
@@ 5.2.1 Dual-Path Search Algorithm (DPSA)
To facilitate the understanding of DPSA, we first introduce some notions related to the BFS tree. The notation $\mathcal{L}_{T_{v^{*}}}$ represents the collection of all leaf nodes of the BFS tree $T_{v^{*}}$. $L_{(v^{*},v_{j})}$ represents the collection of nodes on the shortest path from $v^{*}$ to any node $v_{j}\in\mathcal{L}_{T_{v^{*}}}$ (including $v_{j}$ but not $v^{*}$), i.e., $L_{(v^{*},v_{j})}=\mathcal{SP}(v^{*},v_{j})\backslash v^{*}$. Furthermore, $S_{L_{(v^{*},v_{j})}}$ represents the union of cell-based datasets associated with nodes in the path $L_{(v^{*},v_{j})}$, i.e., $S_{L_{(v^{*},v_{j})}}=\cup_{v_{j}\in L_{(v^{*},v_{j})}}S_{D_{j}}$. $p(S_{L_{(v^{*},v_{j})}})$ represents the price of the path $L_{(v^{*},v_{j})}$, which is the sum of the prices of cell-based datasets corresponding to all nodes on the path $L_{(v^{*},v_{j})}$, i.e., $p(S_{L_{(v^{*},v_{j})}})=\sum_{v_{j}\in L_{(v^{*},v_{j})}}p(S_{D_{j}})$. $\mathcal{S}_{\mathcal{L}_{T_{v^{*}}}}$ represents the collection of $S_{L_{(v^{*},v_{j})}}$ for each $v_{j}\in\mathcal{L}_{T_{v^{*}}}$, i.e., $\mathcal{S}_{\mathcal{L}_{T_{v^{*}}}}=\{S_{L_{(v^{*},v_{j})}},v_{j}\in\mathcal{L}_{T_{v^{*}}}\}$. $\mathcal{R}$ is used to record the union of nodes on the selected path. In the following, we describe the details of DPSA according to the pseudocode of Algorithm 2.
Step 1: Construction of Spatial Dataset Graph. We begin by filtering datasets in the data marketplace $\mathcal{S}_{\mathcal{D}}$ that exceed the allocated budget, resulting in a subset of candidate datasets $\mathcal{S}_{\mathcal{D}}^{\prime}$ (see Lines 2 to 2). We then construct the spatial dataset graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}}^{\prime},\delta}=(\mathcal{V},\mathcal{E}))$ to support the execution of subsequent steps (see Line 2). Specifically, for each node $v_{i}\in\mathcal{V}$, we first compute the cell-based dataset distance (defined in Def. 5) between $v_{i}$ and each node in $\mathcal{V}$. If the distance does not exceed the distance threshold $\delta$, we add an edge to the spatial dataset graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}}^{\prime},\delta}$. The dataset graph construction process continues until all connected dataset pairs are identified.
@@
Step 4. Generating Feasible Solutions.  In the first-round search, the BudgetedGreedy function (see Lines 2 to 2) selects the leaf node $v_{j}$ from $\mathcal{L}_{T_{v^{*}}}$ with the maximum spatial coverage increment $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|$ in each iteration (see Line 2). Then, the function returns the result subset, which is a set of datasets corresponding to all nodes along the selected paths, until the candidate set traversal ends or the remaining budget is exhausted. Similarly, to further prevent the algorithm from falling into a local optimal solution, the function is called again to obtain an alternative feasible solution. This is accomplished by selecting the leaf node $v_{j}$ with the maximum ratio $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|/\Delta_{p}$ in each iteration (see Line 2). Finally, the algorithm returns the better result subset $\mathcal{H}$ from two feasible solutions.
Figure 6: Illustration of DPSA; (a) shows the optimal solution $\mathcal{H}^{*}$ under $B=\$14$; (b) shows the BFS tree $T_{v_{2}}$ rooted at the center $v_{2}$; (c) shows the solution of DPSA $\mathcal{H}=\mathcal{H}^{*}$.
@@
Then the algorithm proceeds with the next iteration, and finds the shortest path $L_{(v_{2},v_{3})}$ that satisfies the budget and maximizes the coverage increment $|S_{L_{(v_{2},v_{3})}}\cap\mathcal{U}|$. Since the node $v_{1}$ is already included in the set $\mathcal{R}$, the cost of adding path $L_{(v_{2},v_{3})}$ is only \$$\$2$. Subsequently, the algorithm continues to iterate and selects $L_{(v_{2},v_{6})}$, leaving a remaining budget of \$$\$2$. At this stage, there is only one candidate path $L_{(v_{2},v_{8})}$ available. However, its price exceeds the remaining budget, resulting in the termination of the iteration. During the second function call, the algorithm iteratively selects paths with the maximum ratio $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|/\Delta_{p}$ ($v_{j}\in\mathcal{L}_{T_{v^{*}}}$) and satisfies all constraints. The shortest paths are added in order of $L_{(v_{2},v_{6})}$, $L_{(v_{2},v_{5})}$, and $L_{(v_{2},v_{3})}$. In this example, both function calls yield an optimal solution, despite the second function call adding paths in a different order compared to the first. The results of DPSA are illustrated in Fig. 6(c).
@@
(1) We first analyze the approximation ratio of the first function call of Algorithm 2. We define $\mathcal{H}_{k^{\prime}}$ as the output after $k^{\prime}$ iterations during the first function call. Additionally, let $\mathcal{H}_{path}^{\prime}$ be the optimal solution of the maximum $k^{\prime}$-coverage problem [18] from $\mathcal{S}_{\mathcal{L}_{T_{v^{*}}}}$. We observe that $\mathcal{H}_{k^{\prime}}$ can be regarded as the output of the greedy algorithm [18] for the maximum $k^{\prime}-$coverage problem. This is because all paths $L_{(v^{*},v_{j})}(v_{j}\in\mathcal{L}_{T_{v^{*}}})$ have the common root node $v^{*}$, which forms a connected subgraph, and the total price is less than the budget $B$. For ease of presentation, we use $|\mathcal{H}_{k^{\prime}}|=|\bigcup_{S\in\mathcal{H}}S|$ to represent the spatial coverage of $\mathcal{H}_{k^{\prime}}$. By Theorem 1 in the maximum coverage problem [18], we have
@@
(2) The second function call is to iteratively select the node with the maximum ratio $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|/\Delta_{p}$. Although it cannot provide any approximation guarantees in theory, but are highly effective in practice, as confirmed by our experiments. In summary, DPSA provides a $\frac{p_{min}}{r\times p_{max}}(1-\frac{p_{max}}{B})(1-\frac{1}{e})$-approximation for the BMCC.

### v3

@@ 3.2 Problem Definition
Suppose we have a marketplace $\mathcal{D}=\{D_{1},\dots,D_{5}\}$, the corresponding dataset prices are {\$$\{\$5,\$6,\$4,\$4,\$2\}$. Fig. 3(a) shows the corresponding cell-based datasets $\mathcal{S}_{\mathcal{D}}$ constructed form $\mathcal{D}$ and a grid $\mathcal{C}_{3}$. We then construct the spatial dataset graph $\mathcal{D}_{\mathcal{S_{D}},\delta}$ with $\delta=2$. For one of the dataset pairs $S_{D_{1}}$ and $S_{D_{2}}$, we can obtain the cell-based dataset distance by calculating the distance between their nearest neighbor cells, as shown in Fig. 3(b). Since the distance $dist(S_{D_{1}},S_{D_{2}})=1<\delta$, $S_{D_{1}}$ and $S_{D_{2}}$ are directly connected. After computing the distance between all pairs of datasets, we can obtain the spatial dataset graph, as shown in Fig. 3(c). A data buyer with a budget of \$15 wants to buy a set of connected datasets. From all feasible solutions of BMCC we can obtain the optimal subset =$\mathcal{H}^{*}=\{S_{D_{1}},S_{D_{2}},S_{D_{4}}\}$.
@@ 4 NP-hardness of BMCC
Given an arbitrary instance of the MCP, we sort the universe $\Omega$ according to the lexicographic order, then create a mapping from the sorted universe $\Omega=\{o_{1},o_{2},\dots,o_{|\Omega|}\}$ to an integer set $\{0,1,\dots,|\Omega|-1\}$. Then we can construct a $2^{\theta}\times 2^{\theta}$ grid in the space such that $2^{\theta}\times 2^{\theta}\geq|\Omega|$. Since each cell in the grid can be represented by a unique integer ID, each element in $\Omega$ can be mapped to a cell whose ID is equal to the integer, as shown in Figs. 4(a) and 4(b). The other four cells with IDs 12, 13, 14, and 15 are not included in $\forall S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$, indicating that no datasets are located in those cells.
Then we set the price of each vertex to 1, the budget to $k$, and $\delta$ to $2^{\theta}\sqrt{2}$ in Definition 7. The set $\mathcal{S}_{\mathcal{D}}$ can be modeled into a complete undirected graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D},\delta}}=(\mathcal{V},\mathcal{E})$: each vertex $v_{i}$ corresponding to each cell-based dataset $S_{D_{i}}\in\mathcal{S}_{\mathcal{D}}$; for each vertex $v_{i}$ in the graph, there is an undirected edge $e_{i,j}$ $(\forall v_{j}\in\mathcal{V}$, $i\neq j)$. This is because $dist(S_{D_{i}},S_{D_{j}})$ must not exceed $\delta$, ensuring the connectivity constraint is always satisfied. Fig. 4(c) shows a spatial dataset graph constructed from the MCP instance, and we can see that the total reduction is performed in polynomial time.
@@ 5 Approximation Algorithms for BMCC
In this section, we focus on solving the BMCC problem (defined in Def. 7). By Theorem 4.1, we know that this problem is NP-hard. Achieving the optimal solution for BMCC is computationally prohibitive. Therefore, we first propose a greedy algorithm, which provides a $\frac{p_{min}}{B}$ approximation guarantee (see Section 5.1). However, its approximation ratio decreases when confronted with a large budget $B$. To address this, we propose an improved greedy algorithm that produces a superior approximation ratio for scenarios involving large $B$ (see Section 5.2).
@@ 5.1 Dual-Search Algorithm (DSA)
In this subsection, we present a Dual-Search Algorithm called DSA. The basic idea of DSA is to perform a two-round heuristic search based on the greedy strategy, the reason for performing the two-round search is primarily the need to balance the three factors of spatial coverage, budget, and spatial connectivity. The goal of the first-round search is to make optimal use of the budget, that is, to maximize the increment of spatial coverage per unit budget. Specifically, the algorithm selects the dataset that maximizes the price-to-coverage ratio (e.g., the ratio of the spatial coverage increment over the price of the dataset) while maintaining the spatial connectivity at each step.
The first-round search ensures efficient use of the budget by considering the price-to-coverage ratio. However, due to the spatial connectivity constraint, it may overlook datasets with larger spatial coverage. For instance, if all datasets have the same price-to-coverage ratio, the algorithm cannot effectively differentiate their cost performance, potentially missing datasets with larger coverage. This results in a locally optimal solution. Therefore, the second-round search iteratively chooses the set that maximizes the marginal gain (e.g., the spatial coverage increment) while maintaining the spatial connectivity.
The second-round search complements the dataset ignored in the first round by maximizing coverage, ensuring maximum spatial coverage of the final result. The optimal selections in the two-round search can result in two feasible solutions, which jointly guarantee the effective approximation ratio of DSA whenever the dataset chooses any pricing function, avoiding local optimal solutions. In the following, we will introduce how DSA works according to the pseudocode of Algorithm 1 and then theoretically prove the approximation ratio of DSA. Afterwards, we will further analyze its time complexity in detail.
@@ 25                   $\mathcal{U}_{2}=\mathcal{U}_{2}\backslash S_{D_{i}}$;
Fig. 5(a) shows a set of cell-based datasets $\mathcal{S}_{\mathcal{D}}=\{S_{D_{1}},\dots,S_{D_{8}}\}$, a distance threshold $\delta=1$ and a budget $B=\$6$. When performing DSA, we first find the candidates $\mathcal{S}_{\mathcal{D}}^{\prime}$. Since the price of each dataset $S_{D}\in\mathcal{S}_{\mathcal{D}}$ is less than $B$, the candidate datasets $\mathcal{S}_{\mathcal{D}}^{\prime}=\mathcal{S}_{\mathcal{D}}$ in this example. The first-round search preferentially selects the dataset with the maximum ratio $|S_{D_{i}}\cap\mathcal{U}_{1}|/{p(S_{D_{i}})}$ within the budget in each iteration. As shown in Fig. 5(b), the algorithm adds $\{v_{8},v_{7},v_{4}\}$ to the result subset $\mathcal{H}_{1}$ one by one until the budget is exhausted, where the spatial coverage of $\mathcal{H}_{1}$ is 8. Next, the second-round search iteratively selects the dataset with the maximum coverage increment $|S_{D_{i}}\cap\mathcal{U}|$ and satisfies spatial connectivity. Thus, we add the set $\{v_{1},v_{5}\}$ into $\mathcal{H}_{2}$, resulting in a spatial coverage of 10 for $\mathcal{H}_{2}$, as shown in Fig. 5(c). Finally, the algorithm returns the result $\mathcal{H}_{2}$, which contains a larger spatial coverage than $\mathcal{H}_{1}$.
@@
Due to the maximum spatial coverage of a set in the result subset not being less than the average spatial coverage of all sets, we have
@@ 38             $\mathcal{H}\leftarrow\mathcal{H}\cup\{S_{D_{i}}\}$;
In the next subsection, we will describe the details of DPSA. Essentially, we try to find a collection of paths with a common node in the spatial dataset graph (defined in Def. 6), where the path is represented as a sequence of nodes connected by edges. This ensures that the subset found is a connected subgraph. However, finding a combination of paths with a common node that maximizes the approximation ratio can be time-consuming (see the time complexity analysis in Section 5.2.1). To improve the efficiency, we propose two accelerate strategies that produce competitive solutions with DPSA in practice while achieving significant speedups in Section 5.2.2.
@@ 5.2.1 Dual-Path Search Algorithm (DPSA)
To facilitate the understanding of DPSA, we first introduce some notions related to the BFS tree. The notation $\mathcal{L}_{T_{v^{*}}}$ represents the collection of all leaf nodes of the BFS tree $T_{v^{*}}$. $L_{(v^{*},v_{j})}$ represents the collection of nodes on the shortest path from $v^{*}$ to any node $v_{j}\in\mathcal{L}_{T_{v^{*}}}$ (including $v_{j}$ but not $v^{*}$), i.e., $L_{(v^{*},v_{j})}=\mathcal{SP}(v^{*},v_{j})\backslash v^{*}$. Furthermore, $S_{L_{(v^{*},v_{j})}}$ represents the union of cell-based datasets associated with nodes in the path $L_{(v^{*},v_{j})}$, i.e., $S_{L_{(v^{*},v_{j})}}=\cup_{v_{j}\in L_{(v^{*},v_{j})}}S_{D_{j}}$. $p(S_{L_{(v^{*},v_{j})}})$ represents the price of the path $L_{(v^{*},v_{j})}$, which is the sum of the prices of cell-based datasets corresponding to all nodes on the path $L_{(v^{*},v_{j})}$, i.e., $p(S_{L_{(v^{*},v_{j})}})=\sum_{v_{j}\in L_{(v^{*},v_{j})}}p(S_{D_{j}})$. $\mathcal{S}_{\mathcal{L}_{T_{v^{*}}}}$ represents the collection of $S_{L_{(v^{*},v_{j})}}$ for each $v_{j}\in\mathcal{L}_{T_{v^{*}}}$, i.e., $\mathcal{S}_{\mathcal{L}_{T_{v^{*}}}}=\{S_{L_{(v^{*},v_{j})}},v_{j}\in\mathcal{L}_{T_{v^{*}}}\}$. $\mathcal{H}$ is used to record the union of nodes on the selected path. In the following, we describe the details of DPSA according to the pseudocode of Algorithm 2.
Step 1: Construction of Spatial Dataset Graph. We begin by filtering datasets in the data marketplace $\mathcal{S}_{\mathcal{D}}$ that exceed the allocated budget, resulting in a subset of candidate datasets $\mathcal{S}_{\mathcal{D}}^{\prime}$ (see Lines 2 to 2). We then construct the spatial dataset graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}}^{\prime},\delta}=(\mathcal{V},\mathcal{E}))$ to support the execution of subsequent steps (see Line 2). Specifically, for each node $v_{i}\in\mathcal{V}$, we first compute the cell-based dataset distance (defined in Def. 4) between $v_{i}$ and each node in $\mathcal{V}$. If the distance does not exceed the distance threshold $\delta$, we add an edge to the spatial dataset graph $\mathcal{G}_{\mathcal{S}_{\mathcal{D}}^{\prime},\delta}$. The dataset graph construction process continues until all connected dataset pairs are identified.
@@
Step 4. Generating Feasible Solutions.  In the first-round search, the BudgetedGreedy function (see Lines 2 to 2) selects the leaf node $v_{j}$ from $\mathcal{L}_{T_{v^{*}}}$ with the maximum ratio $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|/\Delta_{p}$ in each iteration (see Line 2). Then, the function returns the result subset, which is a set of datasets corresponding to all nodes along the selected paths, until the candidate set traversal ends or the remaining budget is exhausted. Similarly, to further prevent the algorithm from falling into a local optimal solution, the function is called again to obtain an alternative feasible solution. This is accomplished by selecting the leaf node $v_{j}$ with the maximum spatial coverage increment $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|$ in each iteration (see Line 2). Finally, the algorithm returns the better result subset $\mathcal{H}$ from two feasible solutions.
Figure 6: Illustration of DPSA, where (a) shows the optimal solution $\mathcal{H}^{*}$ under $B=\$14$; (b) shows the BFS tree $T_{v_{2}}$ rooted at the center $v_{2}$, (c) shows the solution $\mathcal{H}_{1}$ found by the first-round search, and (d) shows the solution $\mathcal{H}_{2}$ found by the second-round search.
@@
To facilitate the understanding of DPSA, we use an example similar to Fig. 5 for illustration. Fig. 6(a) shows the optimal solution of BMCC in $\mathcal{S}_{\mathcal{D}}^{\prime}$ when $B$ is increased to \$$\$14$. Algorithm 2 first performs the pre-processing step, which runs BFS on each node $v$ and calculates its eccentricity $\xi(v)$. Notably, node $v_{2}$ exhibits a minimum eccentricity of 3. Thus, $v_{2}$ is selected as the center and $3$ as the radius $r$. Fig. 6(b) shows the BFS tree $T_{v_{2}}$ rooted at $v_{2}$. All leaf nodes form a subset $\mathcal{L}_{T_{v_{2}}}$=$\{v_{3},v_{5},v_{6},v_{8}\}$. For each node $v_{j}$in $\mathcal{L}_{T_{v_{2}}}$, the union of the cell-based datasets of all nodes on the shortest path $L_{(v_{2},v_{j})}$ forms a new set $\mathcal{S}_{\mathcal{L}_{T_{v_{2}}}}$=$\{S_{L_{(v_{2},v_{3})}},S_{L_{(v_{2},v_{5})}},S_{L_{(v_{2},v_{6})}},S_{L_{(v_{2},v_{8})}}\}$.
@@
During the second function call, DPSA iteratively selects the path with the maximum spatial coverage increment $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|$ ($v_{j}\in\mathcal{L}_{T_{v^{*}}}$) among all available paths. DPSA first adds the node $v_{2}$ to the result subset $\mathcal{H}_{2}$. Next, the spatial coverage increment $|S_{L_{(v_{2},v_{j})}}\cap\mathcal{U}|$ for the four paths from the root node $v_{2}$ to leaf nodes $v_{3},v_{5},v_{8}$ and $v_{6}$ are 9, 10, 8, and 4, respectively. Therefore, DPSA selects $L_{(v_{2},v_{5})}$, which offers the maximum coverage increment $|S_{L_{(v_{2},v_{5})}}\cap\mathcal{U}|$ = 10. After selecting the path $L(v2,v5)$, the remaining budget is calculated as $612-p(S_{D_{1}})-p(S_{D_{5}})=\$12-\$4-\$2=\$$ 6.
The function then proceeds with the next iteration. The coverage increments of the remaining three paths from $v_{2}$ to $v_{3},v_{8}$ and $v_{6}$ are 3, 8, and 4, respectively, with corresponding costs of \$2, \$6, and \$2. Thus, DPSA selects the path $L(v_{2},v_{8})$, which offers the maximum coverage increment 8. At this point, the budget is completely exhausted, and the function terminates, yielding the result set $\mathcal{H}_{2}=\{v_{2},L_{(v_{2},v_{5})},L_{(v_{2},v_{8})}\}$, as shown in Fig. 6(d). Finally, since the spatial coverage $|\mathcal{H}_{1}|=20<|\mathcal{H}_{2}|=21$, the algorithm returns $\mathcal{H}_{2}$, which is identical to the optimal solution $\mathcal{H}^{*}$.
@@
(1) We first analyze the approximation ratio of the second function call of Algorithm 2. We define $\mathcal{H}_{k^{\prime}}$ as the output after $k^{\prime}$ iterations during the first function call. Additionally, let $\mathcal{H}_{path}^{\prime}$ be the optimal solution of the maximum $k^{\prime}$-coverage problem [9] from $\mathcal{S}_{\mathcal{L}_{T_{v^{*}}}}$. We observe that $\mathcal{H}_{k^{\prime}}$ can be regarded as the output of the greedy algorithm [9] for the maximum $k^{\prime}-$coverage problem. This is because all paths $L_{(v^{*},v_{j})}(v_{j}\in\mathcal{L}_{T_{v^{*}}})$ have the common root node $v^{*}$, which forms a connected subgraph, and the total price is less than the budget $B$. For ease of presentation, we use $|\mathcal{H}_{k^{\prime}}|=|\bigcup_{S\in\mathcal{H}}S|$ to represent the spatial coverage of $\mathcal{H}_{k^{\prime}}$. By Theorem 1 in the maximum coverage problem [9], we have
@@
(2) The first function call is to iteratively select the node with the maximum ratio $|S_{L_{(v^{*},v_{j})}}\cap\mathcal{U}|/\Delta_{p}$. Although it cannot provide any approximation guarantees in theory, but are highly effective in practice, as confirmed by our experiments. In summary, DPSA provides a $\frac{p_{min}}{r\times p_{max}}(1-\frac{p_{max}}{B})(1-\frac{1}{e})$-approximation for the BMCC.